OpenStudy (anonymous):

Find the sum. –10 + 1 + (–6) –15 –17 –3 5 @nathan917

OpenStudy (anonymous):

oh that's fine

OpenStudy (anonymous):

@nathan917

OpenStudy (nathan917):

-6 + -10+ 1= -15

OpenStudy (anonymous):

Find the difference. –22 – 5=? –13 –17 –27 27

OpenStudy (nathan917):

-5 + ?= –22 @heyheyhelpme

OpenStudy (anonymous):

the is -27 right?

OpenStudy (nathan917):

-5 + -17 = - 22 So the answer is -17.

OpenStudy (anonymous):

but the question is -22+-5=-27

OpenStudy (nathan917):

Yes you are correct. for some reason im duh-du-duh today i just woke up :P

OpenStudy (anonymous):

sorry if i'm putting stress on you :(

OpenStudy (nathan917):

No. Your fine.

OpenStudy (nathan917):

*slapping myself* That should Do it.

OpenStudy (anonymous):

Find the product. –5(–8) –48 –40 40<<<<<< 35

OpenStudy (nathan917):

I keep getting random messages..

OpenStudy (anonymous):

haha poor thing

OpenStudy (nathan917):

Yes you are correct.

OpenStudy (anonymous):

Find the product. 12 × (–4) × (–5) –32 –240 32 240<<<<<<

OpenStudy (nathan917):

Yes.

OpenStudy (nathan917):

@kelliegirl33 Would you like to join it.?

OpenStudy (anonymous):

Find the quotient. –18 ÷ (–2) (1 point) 9 1/9<<<<<< –9 -1/9

OpenStudy (anonymous):

9

OpenStudy (anonymous):

-18/-2 = 9

OpenStudy (anonymous):

don't you have to start with the ()'s

OpenStudy (anonymous):

@nathan917 @kelliegirl33

OpenStudy (chmvijay):

???

OpenStudy (anonymous):

you don't even need a parenthesis in your last problem.

OpenStudy (nathan917):

Yes.

OpenStudy (anonymous):

@nathan917 is the answer 9

OpenStudy (nathan917):

We are confusing chmvijay lol

OpenStudy (nathan917):

Yes. i would say so.

OpenStudy (anonymous):

cool...we agree

OpenStudy (anonymous):

these are not that hard

OpenStudy (anonymous):

-9 or +9

OpenStudy (nathan917):

Shes in 6th grade @kelliegirl33 This is when she learns them.

OpenStudy (anonymous):

-18/-2 = 9 (positive) A negative divided by a negative equals a psitive

OpenStudy (anonymous):

Once you learn the rules @heyheyhelpme ...it will become easy for you

OpenStudy (anonymous):

Which pair of expressions has the same value? (–2)(–6) and 4 + (–16) (2)(6) and 4 + 16 (–2)(6) and 16 + (–4) (2)(–6) and 4 – 16

OpenStudy (anonymous):

(2)(-6) = 4 - 16 -12 = - 12 answer choice 4

OpenStudy (anonymous):

let me explain something to you....when multiplying or dividing... + x + = + - x - = + - x + = - - divided by - = + - divided by + = -

OpenStudy (nathan917):

I just told i mod off i feel so good :D

OpenStudy (anonymous):

Don't get this confused with addition/subtraction...they are different

OpenStudy (nathan917):

And he left the question :D

OpenStudy (anonymous):

what did you do nathan......

OpenStudy (anonymous):

be careful...you can get suspended that way

OpenStudy (nathan917):

lexiamee... I don't understand mods. Matter in fact a lot of Os member don't also.

OpenStudy (nathan917):

Was not really that bad..

OpenStudy (anonymous):

ooohhh some excitement.....bad boy nathan...lol

OpenStudy (nathan917):

please don't give answers . explain the answer to them . Thats what she/he said.

OpenStudy (nathan917):

i explained it to her..

OpenStudy (anonymous):

@nathan917 is it answer 4

OpenStudy (anonymous):

I occasionally give answers just to help someone out. Some questions you can't really explain. What if the question was "who was the first president of the U.S.?" How do you answer that without giving the answer ?

OpenStudy (nathan917):

Yes. i agree with that.

OpenStudy (nathan917):

Yes the answer is the last one. what would you say keli.

OpenStudy (anonymous):

that is what I said....number 4 (last one)

OpenStudy (anonymous):

okay

OpenStudy (nathan917):

*(2)(-6) = 4 - 16 -12 = - 12 answer choice 4* Ahhh.' But she has to ask me because i have to agree with you. :P

OpenStudy (anonymous):

learn the rules heyheyhelpme....it will make it a whole lot easier

OpenStudy (anonymous):

I'll be right back my Mom needs me

OpenStudy (anonymous):

been there.....moms are important

OpenStudy (nathan917):

Ok take your time.

OpenStudy (anonymous):

okay thanks

OpenStudy (anonymous):

nathan....I am gonna go find some questions to answer....when she comes back she will flag me

OpenStudy (nathan917):

Same.

OpenStudy (anonymous):

@nathan917 @kelliegirl33 I'm back

OpenStudy (anonymous):

@nathan917 @kelliegirl33 I'm back

OpenStudy (anonymous):

:(